Write the equation of the line that passes through point (0, –9) with a slope of 6.
i wrote one but i am not sure that i got this correct can you help? thank you

Answer by dolly(163)

Slope of the line = 6
It passes through (0, -9)
Equation of the line is y - y1 = m(x - x1)
==> y - (-9) = 6(x - 0)
==> y + 9 = 6x is the required equation